

********************************************************************
UARS WEEKLY STATUS REPORT - 21 OCT 2005
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004273 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004274 and 2005099 has been processed and is waiting for archival
at the DAAC.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004259. MEPS data between 2004260 and
2005099 has been processed and is waiting for DAAC archival.
HEPSB data files are version 1. HEPSB data files between 1991274 and
1991366 have been processed and are waiting for DAAC archival.
Processing of all PEM DAAC data files has been suspended until the
DAAC catches up with archiving all processed data files. An error was
discovered which caused problems with DAAC access of PEM Level 2 data.
This was corrected and the DAAC has successfully been able to access SwRI
again. SwRI is waiting for the DAAC to catch up on Level 2 PEM files.
Checks on the PEM data availability through the DAAC have been conducted
by SwRI. These tests were prompted by a user of PEM data who has had trouble
accessing the PEM data in the DAAC. The documentation link to the PEM HEPSA
Level 2 Document still points to the old version file pem_l2_hepsa_v01.txt
instead of the v02 file for hepsa data. Also, there is no MEPS entry for the
MEPS data set. These problems have been reported to help@daac.gsfc.nasa.gov
in mid-October.
Tests of the AXIS1 data file formats have been successful. The AXIS1
format design has been coded. A test file has been generated. A check
program with processing scripts have also been generated. Example binary
read program is complete and verified. These results have been sent to
the project. Creation of the AXIS1 data documents for DAAC data will
begin shortly.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered since updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG is participating in the on-going UARS end of mission
planning activities.
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 08 October 2005 (Orbit 77048, GMT Day 281) through
14 October 2005 (Orbit 77153, GMT Day 287).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
-NONE THIS REPORT PERIOD
The beta angle ranged between the angles of 60.1 degrees to 41.9
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of -2 degrees on 24 October 2005 (DOY 297).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. TMON 45 and 46 are
maintaining the relay operations for battery 2. Battery 2 and Battery
3 are currently maintaining greater than 23.6 V end-of-night (EON)
load bus voltage. Battery temperatures are stable with a temperature
delta between Battery 2 and Battery 3 of 10.9 to 10.2 degrees C.
Additional battery performance data for each battery on 07 October
2005 and 14 October 2005 are provided in Appendix D.
The clock error ranged from 10.6 to -11.1 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 06
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 04
Five (5) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Four (4) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continues to closely monitor the condition of
battery 2 and battery 3. Battery 2 and 3 have not been sharing load
well since the cell failure on battery 2. Battery 3 now carries
almost 90% of the power load while battery 2 provides the remaining
10%. NASA power branch has been providing daily support in monitoring
the UARS battery condition. The FOT has been authorized to disable
all platform and instrument heaters on the SSPP. This will be done at
the earliest available opportunity. The FOT also changed VT levels
from 7 to 6 because of the approaching high beta. The FOT will
revert back to VT-7 at a future date based on power needs.
The FOT successfully completed the one planned orbit lowering burn
this week. The orbit lowering burn on 12 October 2005 went as
expected. The burn started as planned at 285/1509 and ran for 18
minutes. The calculated duty cycle for the burn was 77.7%. FDF
calculated that perigee was lowered by 30.4 km at the completion of
this burn.
MEETINGS
The FOT attended a combined post burn de-brief and CAM meeting on 11
October 2005. The FOT and FDF presented their analysis of the 18-
minute burn that occurred on 06 October 2005. The FOT re-sequenced a
few commands in the burn script in order to minimize the time needed
before the burn start time. The orbit lowering burn script and
procedures for the burn on 12 October 2005 were reviewed and approved
during this meeting. FDF also presented their predicted post burn
position and velocity based on the engineering calibration burns done.
The FOT now only needs 3 minutes of commanding time from the start of
the pass to be able to fire the thrusters, and 5 minutes to safe the
propulsion module before LOS of the burn pass.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 24 October 2005 (DOY 297).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
UARS has been given permission to execute the scheduled 18-minute
burns on 18 October 2005 and 20 October 2005. These two burns are the
last two that will be executed during this solar beta cycle. UARS
will yaw around to the forward direction on 24 October 2005 (DOY 297).
The FOT will be able to resume orbit-lowering burns after the forward
to reverse yaw maneuver on 23 November 2005.
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 14 October 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
77048 281/0209 05-363 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
77077 283/0213 05-364 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
77116 285/1220 05-365 PM ORBIT ADJUST
77132 286/1425 05-366 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
77149 287/1802 05-369 OBC/EPHEM WINOPEN, EPH LOAD
77149 287/1802 05-368 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
77152 287/2219 05-369 OBC/EPHEM WINDEF
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 14 October 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 14 October 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
77048 281/0212 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
77077 283/0215 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
77132 286/1427 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
77149 287/1804 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 14 October 2005
07 October 2005 - GMT Day 280 Beta = 61.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -1.79/ -3.05 +00.0/+00.0 04.6 0.00 0.0
2 -700.0/-700.0 -1.17/ -2.42 +09.6/-05.6 25.8 1.08 2.3
3 324.8/ 056.0 +9.73/ +7.74 +37.6/-27.6 25.8 1.05 25.8
14 October 2005 - GMT Day 287 Beta = 41.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -1.48/ -2.42 +00.0/+00.0 04.6 0.00 0.0
2 -700.0/-700.0 -0.54/ -1.79 +13.6/-06.8 25.4 1.05 3.3
3 011.2/-134.4 +10.40/ +8.40 +44.4/-25.6 25.4 1.04 25.8
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 14 October 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-047 FOT 77051 281/0712 30832 Generic Late Acquisition
05-049 FOT 77094 284/0157 41076 WSGT unable to support
05-050 FOT 77100 284/1213 41080 Neg Acquisition
05-051 FOT 77101 284/1345 41080 Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-048 FOT 76997 277/1741 41006 Post-Burn Late Acq (05:04)
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-042 MPS 76339 233/2347 N/A BATTERY 2 DIFFV -700mV
05-047 FOT See List Oct 05 30832 Generic Late Acquisition
05-049 FOT 77094 284/0157 41076 WSGT unable to support
05-050 FOT 77100 284/1213 41080 Neg Acquisition
05-051 FOT 77101 284/1345 41080 Late Acquisition